# SynthesizRR: Generating Diverse Datasets with Retrieval Augmentation

### v1

@@ Abstract
Large language models (LLMs) are versatile and can address many tasks, but for computational efficiency, it is often desirable to distill their capabilities into smaller student models. One way to do this for classification tasks is via dataset synthesis, which can be accomplished by generating examples of each label from the LLM. Prior approaches to synthesis use few-shot prompting, which relies on the LLM’s parametric knowledge to generate usable examples. However, this leads to issues of repetition, bias towards popular entities, and stylistic differences from human text. In this work, we propose Synthesize by Retrieval and Refinement (SynthesizRR), which uses retrieval augmentation to introduce variety into the dataset synthesis process: as retrieved passages vary, the LLM is “seeded” with different content to generate its examples. We empirically study the synthesis of six datasets, covering topic classification, sentiment analysis, tone detection, and humor, requiring complex synthesis strategies. We find SynthesizRR greatly improves lexical and semantic diversity, similarity to human-written text, and distillation performance, when compared to standard 32-shot prompting and six baseline approaches.
@@ 1 Introduction
Figure 1: Synthetic examples from few-shot generation (middle) and SynthesizRR (bottom). Few-shot generation leverages a frozen LLM’s parametric knowledge as the source of content in the generated examples. SynthesizRR incorporates a distinct content sourcing step by retrieving documents from a corpus. For the task of detecting political bias, a news article is retrieved and transformed in tone. The resulting synthesis procedure yields diverse examples which more closely match human-written examples (top).
@@
Figure 2: Abstract depiction of the SynthesizRR pipeline. In the content sourcing stage, we retrieve $K$ unique document $\{r_{1},\dots,r_{K}\}$ from a large corpus for each in-context covariate $x_{\text{ICL}}$. The task-inversion stage of synthesis uses a parameterized context refinement prompt $\mathcal{P}_{\tau}$, which takes parameters $R_{inv}$ (inversion instruction), $r_{k}$ (a retrieved document), and $\mathcal{V}(y_{\text{ICL}})$ (the verbalized target label). A generalist teacher LLM autoregressively generates a synthetic covariate. Each in-context example thus produces $K$ unique synthetic examples $\{\tilde{x}_{1},\dots,\tilde{x}_{K}\}$, which we include in the dataset with target $y_{\text{ICL}}$.
@@
Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023; Bubeck et al., 2023), LLaMa (Touvron et al., 2023) and Claude (Bai et al., 2022) are powerful generalist models, capable of solving multiple tasks without parameter tuning via zero-shot or few-shot prompting. In comparison, previous approaches fine-tuned variants of BERT (Devlin et al., 2019) on task-specific demonstrations, producing specialist models. These smaller specialist models are more economical at inference time, but typically require at least thousands of examples to train.
@@
Recent work has sought to avoid this reliance on annotation by fine-tuning specialist models on synthetic datasets via teacher-student distillation (West et al., 2022). While this has been employed in the context of classification (Yu et al., 2023a; Ye et al., 2022a, b), human-preference alignment (Lee et al., 2023; Bai et al., 2022), language understanding (Meng et al., 2022; Schick and Schütze, 2021), and even tabular data (Borisov et al., 2022), synthetic data has limitations. Yu et al. (2023a) examine the data-generation process and discover that naive prompts generate text inputs that are low in diversity and regurgitate the biases of the teacher LLMs, which proliferate into the student model.
Figure 1 illustrates this: few-shot generation (Brown et al., 2020) is a common approach that uses an LLM to perform a multitude of tasks without modifying its parameters. Prior work has used few-shot generation to synthesize a training dataset for fine-tuning a smaller model (Ye et al., 2022a, b; Yehudai et al., 2024a). We refer to this type of dataset synthesis as FewGen.
@@
FewGen relies on a generalist LLM’s ability to perform task inversion. For example, in Figure 1, we illustrate the task of detecting politically-biased articles. By using the prompt “Write a news article using harsh political language, a mocking tone and toxic commentary” and a few in-context learning examples, sampling continuations from an LLM is often sufficient to generate plausible news articles in the biased style we seek to detect. However, when thousands of completions are sampled from this prompt to construct a large dataset for distilling a specialist model, we observe repetition, bias towards popular entities, and stylistic differences from human-written texts. If the data is low diversity, a trained specialist model may not learn the task well.
@@
In this work, we seek to alleviate the lack of diversity in synthetic data. We make the observation that dataset synthesis may be decomposed as two distinct LLM competencies: content sourcing, where the LLM obtains relevant information for the task, and task inversion, where the LLM generates a synthetic input example. Prior work has focused mainly on task inversion, i.e., generating good covariates using a target-conditioned prompt, while implicitly using the LLM’s parametric memory for content sourcing. In contrast, we investigate the importance of an explicit content sourcing stage.
@@
We propose Synthesize by Retrieval and Refinement (SynthesizRR [^1] Pronounced “synthesizer”.), a procedure by which a large corpus acts as the guiding signal for the content sourcing stage. In the initial step, we repurpose each in-context learning input as a retrieval query. We use this to retrieve several documents from a large domain-specific corpus. We then use a generalist LLM for task inversion on each retrieved document. As a unique retrieved document is used in each prompt, the synthesis procedure yields more diverse examples, containing a wider range of real-world entities and assertions.
@@
We benchmark SynthesizRR against FewGen on six text classification tasks, selected carefully to measure a variety of different styles of dataset synthesis. Experiments on an in-domain corpus (§5) reveal that SynthesizRR significantly surpasses FewGen in diversity and resemblance to human-authored texts, even though both procedures utilize the same frozen LLM. In §6, we see that student classifiers fine-tuned on SynthesizRR-generated data perform better than those fine-tuned on FewGen. We analyze the contributing factors in the student’s training dynamics, finding that our approach generates examples with superior class-separability, particularly for multiclass tasks.
Finally, in §7, we compare SynthesizRR to six prior approaches for synthesis of classification datasets: AttrPrompt Yu et al. (2023a), ReGen Yu et al. (2023b), SunGen Gao et al. (2023), ZeroGen Ye et al. (2022a), ProGen Ye et al. (2022b), and S3 Wang et al. (2023a). In general, we find that SynthesizRR gives higher diversity data, and in most cases, leads to higher student accuracy as a result.
Our contributions are as follows: (1) we propose a new method for integrating retrieval into example synthesis for teacher-student distillation; (2) we perform a detailed analysis comparing our method with existing approaches across both intrinsic data diversity metrics and downstream task accuracy.
@@ 2 Background and Task setup
The dataset generation tasks we focus on in this paper are in the domain of text classification. Denote an example as consisting of an input $x$, typically a natural language string, and output $y\in\mathcal{Y}$ for an output space $\mathcal{Y}$ of $C$ classes.
We assume access to a seed set of examples $\mathcal{D}_{\textsc{Seed}}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ which is large enough to cover the classes but small enough that a user could feasibly gather or construct these examples manually. In §5, we use the in-context learning set as $\mathcal{D}_{\textsc{Seed}}$.
Our assumption is that $\mathcal{D}_{\textsc{Seed}}$ is not large enough to train an effective student, and instead a larger synthetic dataset $\mathbin{\mathbin{\mathcal{D}_{\textsc{Synth}}}=\left\{(\mathbin{\tilde{x}^{i}},\mathbin{y^{i}})\right\}^{m}_{i=1}}$ ($m>>n$) will be needed. We repeatedly prompt a teacher language model $\mathcal{M}_{\text{LM}}$ to generate synthetic covariates $\tilde{x}$ given corresponding labels $y$ to produce $\mathbin{\mathcal{D}_{\textsc{Synth}}}$, a process we refer to as task inversion. A student language model $\mathcal{M}_{\text{S}}$ (e.g., a BERT-style pre-trained encoder model (Devlin et al., 2019)) is trained on $\mathbin{\mathcal{D}_{\textsc{Synth}}}$. We denote the student’s task (predicting $y$ from $x$) by $\tau$ and the teacher’s task (generating $x$ given $y$) by $\tau_{inv}$.
SynthesizRR aims to address the lack of diversity by leveraging retrieval from a large corpus $\mathcal{R}$ for the content sourcing step. Figure 2 illustrates how our method can be used to generate appropriate data. First, we start by executing a retrieval step based on the examples $\mathcal{D}_{\textsc{Seed}}$. Our assumption here is that we can find similar examples in the corpus, which are domain-similar enough to $\mathcal{D}_{\textsc{Seed}}$ to be usable as training examples. A context refinement prompt is then constructed to perform task inversion on each retrieved document. The LLM’s context window thus sees a unique and grounded prompt for each example it generates. This overcomes the need to memorize a corpus worth of information in the limited parameters of the teacher LLM.
We assume the existence of corpus $\mathcal{R}$, with each document containing information relevant to the task. However, the documents do not need to be unlabeled examples for our task; they can be more distantly related and still provide useful information for generating examples. In §5 we successfully generate customer reviews and questions about products, while retrieving from a corpus where documents are product titles and descriptions. Appendix C looks into how the corpus influences synthesis. Another reason raw retrieval documents are unsuitable for direct use as examples is due to the chunking process in retrieval, which can introduce coreference failures, among other problems (Mialon et al., 2023)).
Due to the difficulty of the task inversion step as a result of this mismatch between retrieved documents and test examples, we limit our investigation to chat-tuned teacher LLMs which have been shown to demonstrate strong instruction-following capabilities (Ouyang et al., 2022; Touvron et al., 2023; Bai et al., 2022).
Algorithm 1 SynthesizRR Retr-ICL
@@
     $\mathcal{D}_{\textsc{Retr}}\leftarrow\mathcal{D}_{\textsc{Retr}}\cup\{(x,y,[r_{1},\ldots,r_{K}])\}$
@@
for $(x,y,[r_{1},\ldots,r_{K}])\in\mathcal{D}_{\textsc{Retr}}$ do
     for $r_{k}\in[r_{1},\ldots,r_{K}]$ do
@@
for $(x,y,[r_{1},\ldots,r_{K}])\in\mathcal{D}_{\textsc{Retr}}$ do
     for $r_{k}\in[r_{1},\ldots,r_{K}]$ do
@@ 3 Method
Algorithm 1 shows our dataset generation method.
Step 1. Retrieval using $\mathcal{M}_{\mathrm{ret}}$: SynthesizRR uses each in-context demonstration $(x_{{}_{\text{ICL}}},y_{{}_{\text{ICL}}})$ as a query for information retrieval, in addition to its role as an in-context demonstration for task inversion as discussed subsequently. When viewed as queries, these demonstrations lead to an expansion as we retrieve $K$ documents $\{r_{1},\dots,r_{K}\}$ of decreasing similarity for each in-context input $x_{{}_{\text{ICL}}}$. We rank contexts using retrieval model $\mathcal{M}_{\mathrm{ret}}$ and retain contexts with cosine similarity in (0.4, 0.9), where the upper limit helps avoid duplicates by excluding retrieved contexts if they are too close to the ICL example. We thus append each triplet $(x_{{}_{\text{ICL}}},y_{{}_{\text{ICL}}},[r_{1},\dots,r_{K}])$ to the set $\mathcal{D}_{\textsc{Retr}}$.
Step 2. Retrieval-augmented task inversion: The retrieved documents have lexical and semantic overlap with the query, but are not necessarily suitable examples for the final dataset. To accommodate this, we use a prompt to perform task-inversion using content sourced from each retrieved document, which we refer to as contextual refinement.
The minimum elements of the task inversion prompt $\mathcal{P_{\tau}}$ are the instruction $\mathcal{I}_{inv}$ and target $y$. We use a verbalizer function (Schick and Schütze, 2021; van de Kar et al., 2022) that provides a unique and descriptive representation of each label, i.e. $\mathcal{V}:\{y_{1},\dots,y_{C}\}\rightarrow\{v_{1},\dots,v_{C}\}$. We follow prior work on classification-based task inversion (Schick and Schütze, 2021; Ye et al., 2022a, b; Yu et al., 2023b; Gao et al., 2023) and use descriptive verbalizations to induce label-separability in the final dataset.
FewGen uses the standard causal language modeling objective to induce next-token probabilities from $\mathcal{M}_{\text{LM}}$. Nucleus sampling Holtzman et al. (2019) is used to autoregressively sample next tokens until the <eos> token is generated. This becomes synthetic example $\mathbin{\tilde{x}}^{i}$.
@@
For each label $y$, we sample $m/C$ times to generate $\mathcal{D}_{\textsc{Synth}}=\bigcup\limits_{y\in\mathcal{Y}}{}\bigcup\limits_{i\in\{1..m/C\}}\left\{(\mathbin{\tilde{x}^{i}},y)\right\}$.
For SynthesizRR, the task inversion prompt is at minimum composed of the contextual refinement instruction $\mathcal{R}_{inv}$ alongside the verbalized target from the query example $\mathcal{V}(y_{{}_{ICL}})$. The LLM’s context window thus sees a unique and grounded prompt when auto-regressively generating each synthetic input $\mathbin{\tilde{x}}^{i}$:
@@
for all documents $\text{for all documents }r_{k}\in\{r_{1},\cdots r_{K}\}$
We continue to use nucleus sampling to get diverse generations. Each in-context example thus produces $K$ unique synthetic examples $\{\tilde{x}_{1},\dots,\tilde{x}_{K}\}$, where we refer to $K$ as the “expansion factor”.
Step 3. Student distillation: The student is fine-tuned on $\mathbin{\mathcal{D}_{\textsc{Synth}}}$ by passing the BERT [CLS] token embedding through a feedforward layer. This produces a probability distribution over the labelspace $C$. We optimize the cross-entropy loss of the true label $y$ (which was used in the task-inversion prompt, thus is known to us) when the student model is passed the generated example $\mathbin{\tilde{x}}$. As we derive $\mathbin{\tilde{x}}$ from a teacher LLM, this can be considered a form of symbolic knowledge distillation (West et al., 2022) without the need for a critic model.
Unlike prior work, we make no changes to the student model to aid the out-of-distribution performance and focus solely on improvements to the dataset synthesis procedure.
In-context learning in SynthesizRR. Contextual refinement may also benefit from in-context demonstrations. In this case, we face a challenge in constructing an in-context learning set which accurately captures our context refinement task $r_{k}^{i}\rightarrow\mathbin{\tilde{x}}^{i}$. A basic exploration using handpicked demonstrations reveals that in-context learning primarily influences the format, rather than content, of synthetic examples. This finding is also supported by the literature on in-context learning (Min et al., 2022). We thus develop two approaches to in-context learning (ICL).
1. Retr-ICL: we use retrieval to construct a set of ICL examples $\mathcal{D}_{\textsc{ICL}}$, such that each ICL example looks similar to the SynthesizRR task-inversion prompt. We consider top-1 and top-2 retrieved results from dense retriever $\mathcal{M}_{\mathrm{ret}}$, and use a cosine-similarity criterion $s_{\alpha}\leq\cos(x_{{}_{\text{ICL}}},r_{k})\leq s_{\beta}$ to determine whether there is a possible match between the retrieved document $r_{k}$ and the ICL covariate $x_{{}_{\text{ICL}}}$. Within a massive corpus, there are typically sufficient semantically-close retrieved results to form an in-context pair ($r_{k},x_{{}_{\text{ICL}}}$). When used with an appropriate prompt template from Appendix E these ICL examples demonstrate the required format and approximately match in content.
2. No-Retr-ICL: a baseline method, where we use retrieval for content sourcing, but not for in-context learning. In No-Retr-ICL, for each generation we select 32 ICL examples at random from $\mathcal{D}_{\textsc{Seed}}$. Each ICL example is preceded by a prefix such as “News Article:” or “Product details:”, but without any instruction. After these ICL examples, we append the retrieved document $r_{k}$ and context refinement instruction $R_{inv}$ to produce the final prompt. The resulting in-context learning prompt is very similar to that used by FewGen, but also uses content-sourcing by including $r_{k}$ and $R_{inv}$. This baseline highlights the importance of using $\mathcal{D}_{\textsc{ICL}}$ in the Retr-ICL approach.
Dataset Class Train, Test In-domain corpus Task inversion
@@
ToI Headlines 10 $52\text{k},10\text{k}$ RN/Ind Easy
Category 23 $30\text{k}^{*},2.4\text{k}$ Products Medium
@@
Polarity 2 $72\text{k}^{*},7.2\text{k}^{*}$ Products Medium
Table 1: Dataset statistics and our estimate of task inversion complexity. Downsampled for convenience.

### v3

@@ Abstract
It is often desirable to distill the capabilities of large language models (LLMs) into smaller student models due to compute and memory constraints. One way to do this for classification tasks is via dataset synthesis, which can be accomplished by generating examples of each label from the LLM. Prior approaches to synthesis use few-shot prompting, which relies on the LLM’s parametric knowledge to generate usable examples. However, this leads to issues of repetition, bias towards popular entities, and stylistic differences from human text. In this work, we propose Synthesize by Retrieval and Refinement (SynthesizRR), which uses retrieval augmentation to introduce variety into the dataset synthesis process: as retrieved passages vary, the LLM is “seeded” with different content to generate its examples. We empirically study the synthesis of six datasets, covering topic classification, sentiment analysis, tone detection, and humor, requiring complex synthesis strategies. We find that SynthesizRR [^1] https://github.com/amazon-science/synthesizrr greatly improves lexical and semantic diversity, similarity to human-written text, and distillation performance, when compared to 32-shot prompting and four prior approaches.
@@ 1 Introduction
Figure 1: Synthetic examples from few-shot generation (middle) and SynthesizRR (bottom). Our approach incorporates a content sourcing step which retrieves documents from a corpus: for the task of detecting political bias, a news article is retrieved and the teacher LLM is prompted to produce a biased version. The resulting synthesis procedure yields diverse examples which more closely match human-written examples.
@@
Figure 2: Abstract depiction of the SynthesizRR procedure. In the content sourcing stage, we retrieve $K$ unique document $\{r_{1},\dots,r_{K}\}$ from a large corpus for each in-context covariate $x_{\text{ICL}}$. The task-inversion stage of synthesis uses a parameterized context refinement prompt $\mathcal{P}_{\tau}$, which takes parameters $R_{inv}$ (inversion instruction), $r_{k}$ (a retrieved document), and $\mathcal{V}(y_{\text{ICL}})$ (the verbalized target label). A generalist teacher LLM autoregressively generates a synthetic covariate. Each in-context example thus produces $K$ unique synthetic examples $\{\tilde{x}_{1},\dots,\tilde{x}_{K}\}$, which we include in the dataset with target $y_{\text{ICL}}$.
@@
Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023; Bubeck et al., 2023), LLaMa (Touvron et al., 2023b) and Claude (Bai et al., 2022) are versatile generalist models, capable of solving multiple tasks without parameter tuning via zero-shot or few-shot prompting. In comparison, previous approaches fine-tuned variants of BERT (Devlin et al., 2019) on task-specific demonstrations, producing specialist models. These smaller specialist models are more economical at inference time, but require at least thousands of examples to train.
@@
Recent work has sought to avoid this reliance on manually created examples by fine-tuning specialist models on synthetic datasets via teacher-student distillation (West et al., 2022). This has applications in classification (Yu et al., 2023a; Ye et al., 2022a, b), human-preference alignment (Lee et al., 2023; Bai et al., 2022), language understanding (Meng et al., 2022; Schick and Schütze, 2021), and even tabular data (Borisov et al., 2022). However, synthetic data has limitations. As Yu et al. (2023a) note, naive prompts generate texts with limited diversity and reflecting biases of the teacher LLMs.
@@
Figure 1 illustrates the few-shot synthesis approach (Ye et al., 2022a, b; Yehudai et al., 2024a), which we refer to as FewGen, for the task of detecting politically-biased articles. With a suitable prompt and in-context examples, sampling continuations from an LLM generates plausible news in the biased style we seek to detect. However, as thousands of completions are sampled from a fixed prompt, we observe repetition, bias towards popular entities, and stylistic differences from human-written texts. Specialist models distilled from such low diversity datasets may not learn the task well.
@@
In this work, we seek to alleviate the lack of diversity in synthetic data. We suggest that dataset synthesis may be decomposed as two distinct LLM competencies: content sourcing, where the LLM obtains relevant information for the task, and task inversion, where the LLM generates a synthetic input using a target-conditioned prompt. Prior work has focused mainly on task inversion, while implicitly using the LLM’s parametric memory for content sourcing. In contrast, we investigate the importance of an explicit content sourcing stage.
@@
We propose Synthesize by Retrieval and Refinement (SynthesizRR), an example synthesis procedure guided by a retrieval corpus. In the content sourcing step, we use in-context learning covariates as retrieval queries to extract dozens of documents per query from a domain-specific corpus. Subsequently, a generalist LLM performs task inversion on each retrieved document. As each prompt uses a unique retrieved document, our synthesis procedure generates diverse examples, enriched with a broad spectrum of real-world entities and assertions.
@@
We benchmark SynthesizRR against FewGen on six text classification tasks, selected carefully to measure a variety of different styles of dataset synthesis. Our experiments (§5) reveal that SynthesizRR significantly surpasses FewGen in diversity and resemblance to human-authored texts, even though both procedures utilize the same frozen LLM. In §6, we see that student classifiers fine-tuned on SynthesizRR-generated data perform better than those fine-tuned on FewGen. Finally, in §7, we compare SynthesizRR to four state of the art approaches for synthesis of classification datasets, and find SynthesizRR gives higher diversity datasets, better matching human-written instances, and leads to higher student accuracy in most cases.
Our contributions are as follows: (1) we propose a new method of example synthesis for teacher-student distillation, which grounds the task inversion step using a retrieval corpus; (2) we introduce the SynthesizRR RetrICL algorithm to create a realistic in-context learning set for our method; (3) we empirically analyze the synthesis of six challenging classification tasks, comparing our method’s textual diversity and similarity and downstream task accuracy to existing approaches; (4) we pinpoint factors affecting the quality of our synthetic datasets by varying the amount of supervised data, corpus relevance to task, number of in-context examples, and sparse vs. dense retrieval.
@@ 2 Background and Task setup
In this paper, we focus on generating datasets for challenging text classification tasks. Denote an example as consisting of input text $x$ and output $y\in\mathcal{Y}$ for output space $\mathcal{Y}$ of $C$ classes. Our goal is to produce a synthetic dataset $\mathbin{\mathbin{\mathcal{D}_{\textsc{Synth}}}=\left\{(\mathbin{\tilde{x}^{i}},\mathbin{y^{i}})\right\}^{m}_{i=1}}$ and train a specialist language model $\mathcal{M}_{\text{S}}$ (e.g. a BERT-style pre-trained model (Devlin et al., 2019)). We create $\mathbin{\mathcal{D}_{\textsc{Synth}}}$ via task inversion: repeatedly prompting a teacher language model $\mathcal{M}_{\text{LM}}$ to generate synthetic covariates $\tilde{x}$ given corresponding labels $y$. We denote the student’s task (predicting $y$ from $x$) as $\tau$ and the teacher’s task (generating $x$ given $y$) as $\tau_{inv}$.
SynthesizRR aims to address the lack of diversity by leveraging retrieval during the content sourcing step. We assume the existence of a corpus $\mathcal{R}$ where each document may hold task-relevant information. However, documents need not originate from the same distribution as our task covariates; even distantly related documents can yield valuable synthetic examples. For instance, we shows that we can successfully generate reviews and humorous questions from a corpus of product descriptions. We also assume access to a seed set of examples $\mathcal{D}_{\textsc{Seed}}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ which is sufficiently large to represent the classes but small enough to be manually compiled by a user in a few hours; in experiments, we use the in-context learning set as $\mathcal{D}_{\textsc{Seed}}$. Importantly, we assume the seed set is insufficient to train an effective student, and a larger $\mathbin{\mathcal{D}_{\textsc{Synth}}}$ ($m>>n$) is needed.
Figure 2 illustrates our method for generating distributionally similar covariates. Initially, we retrieve documents based on the examples in $\mathcal{D}_{\textsc{Seed}}$, assuming that the corpus contains sufficient domain-similar documents. We then construct a context refinement instruction to perform task inversion on each retrieved document. This approach provides the LLM with a unique and grounded prompt for each generated example, thereby circumventing the need for the teacher LLM to memorize extensive corpus data within its limited parameters. Task inversion may be challenging due to the mismatch between retrieved documents and test examples; to overcome this, we limit our investigation to teacher LLMs demonstrating strong instruction-following capabilities (Ouyang et al., 2022; Touvron et al., 2023b; Bai et al., 2022).
Algorithm 1 SynthesizRR RetrICL
@@
     $\Gamma_{K}\leftarrow[r_{1},\ldots,r_{K}]$
     $\mathcal{D}_{\textsc{Retr}}\leftarrow\mathcal{D}_{\textsc{Retr}}\cup\{(x,y,\Gamma_{K})\}$
@@
for $(x,y,\Gamma_{K})\in\mathcal{D}_{\textsc{Retr}}$ do
     for $r_{k}\in\Gamma_{K}$ do
@@
for $(x,y,\Gamma_{K})\in\mathcal{D}_{\textsc{Retr}}$ do
     for $r_{k}\in\Gamma_{K}$ do
@@ 3 Method
Algorithm 1 shows our dataset generation method. We distill a student model in these steps:
Step 1. Content sourcing using retrieval: SynthesizRR uses each in-context covariate $x_{{}_{\text{ICL}}}$ as a query for information retrieval, in addition to its subsequently role during in-context learning. For each query, we retrieve $K$ documents $\Gamma_{K}=[r_{1},\dots,r_{K}]$ of progressively decreasing cosing similarity using the dense retriever $\mathcal{M}\mathrm{ret}$. We retain documents with cosine similarity in (0.4, 0.9), to ensure minimum similarity while excluding overly similar documents as potential duplicates of $x_{{}_{\text{ICL}}}$. Each resulting triplet $(x_{{}_{\text{ICL}}},y_{{}_{\text{ICL}}},\Gamma_{K})$ is appended to set $\mathcal{D}_{\textsc{Retr}}$.
Step 2. In-context set construction: The subsequent task inversion step also benefits from in-context demonstrations, but it is challenging to construct demonstrations which effectively captures our context refinement task $r_{k}^{i}\rightarrow\mathbin{\tilde{x}}^{i}$. We explored two approaches to in-context learning.
1. RetrICL: we use retrieval to construct a set of ICL examples $\mathcal{D}_{\textsc{ICL}}$, such that each ICL example mirrors the format of our task-inversion prompts. We select top-1 and top-2 retrieved results from the densely retrieved results, and use a cosine-similarity criterion $s_{\alpha}\leq\cos(x_{{}_{\text{ICL}}},r_{k})\leq s_{\beta}$ to asses the potential match between the retrieved document $r_{k}$ and $x_{{}_{\text{ICL}}}$. Although the in-context pair may not match exactly, they demonstrate the required format as per Appendix G.
2. Non-RetrICL: a baseline method, which uses retrieval for content sourcing, but not for in-context learning. For each generation we select $N=32$ ICL examples at random from $\mathcal{D}_{\textsc{Seed}}$. Each example is appended with a prefix like “News Article:” or “Product details:” but we do not add the context refinement instruction. After the ICL examples, we append the retrieved document $r_{k}$ and context refinement instruction $R_{inv}$ to form the final prompt. This format closely mirrors the in-context learning prompt used by FewGen, but also incorporates content-sourcing elements $r_{k}$ and $R_{inv}$. This baseline highlights the value added by constructing $\mathcal{D}_{\textsc{ICL}}$ in the RetrICL approach.
Step 3. Task inversion using context refinement: The minimum elements of a task inversion prompt $\mathcal{P_{\tau}}$ are the context refinement instruction $\mathcal{I}_{inv}$ and target $y$. We use a verbalizer function $\mathcal{V}$ (Schick and Schütze, 2021; van de Kar et al., 2022) to provide a unique text representation of each label, i.e. $\mathcal{V}:\mathcal{Y}\rightarrow\{v_{1},\dots,v_{C}\}$. We follow prior work on classification-based task inversion (Schick and Schütze, 2021; Ye et al., 2022a, b; Yu et al., 2023b; Gao et al., 2023) and use descriptive verbalizations to induce label-separability in the final dataset.
FewGen uses the standard causal language modeling objective to induce next-token probabilities from teacher LLM, $\mathcal{M}_{\text{LM}}$. Nucleus sampling Holtzman et al. (2019) is used to autoregressively sample next tokens until the <eos> token is generated. This becomes synthetic example $\mathbin{\tilde{x}}^{i}$.
@@
For each label $y$, we fix this prompt and sample $m/C$ times to generate the synthetic dataset.
In SynthesizRR, we create the synthetic dataset from each triplet in $\mathcal{D}_{\textsc{Retr}}$. The retrieved documents $\Gamma_{K}=[r_{1},\dots,r_{K}]$ have lexical and semantic overlap with the query $x_{{}_{\text{ICL}}}$. However, corpus documents may be distributionally dissimilar from real task covariates, due to the nature of documents or chunking process (Mialon et al., 2023). To address this, we use $\mathcal{M}_{\text{LM}}$ to perform task inversion from the content of each retrieved document, a process we refer to as contextual refinement. $\mathcal{P_{\tau}}$ is thus composed from the contextual refinement instruction $\mathcal{R}_{inv}$, each document $r_{k}\in\Gamma_{K}$, and the verbalized target for the query, i.e. $\mathcal{V}(y_{{}_{ICL}})$. The LLM’s context window thus sees a unique and grounded prompt when auto-regressively generating each synthetic input $\mathbin{\tilde{x}}^{i}$:
@@
for all documents $r_{k}\in\Gamma_{K}$. We continue to use nucleus sampling to get diverse generations. Each original in-context example thus produces $K$ unique synthetic examples $\{\tilde{x}_{1},\dots,\tilde{x}_{K}\}$; we call $K$ the “expansion factor”. To promote adherence to $\mathcal{R}_{inv}$, we sample pairs from $\mathcal{D}_{\textsc{ICL}}$ to create in-context examples following the same format. Our final dataset is constructed as:
$\mathcal{D}_{\textsc{Synth}}=\bigcup\limits_{(x,y,\Gamma_{K})\in\mathcal{D}_{\textsc{Retr}}}{}\bigcup\limits_{r_{k}\in\Gamma_{K}}\left\{(\mathbin{\tilde{x}^{i}},y)\right\}$.
Step 4. Student distillation: The student is fine-tuned on $\mathbin{\mathcal{D}_{\textsc{Synth}}}$ by passing the BERT [CLS] token embedding of $\mathbin{\tilde{x}}$ through a feedforward layer. This produces a probability distribution over the label space $C$. We optimize the cross-entropy loss of the true label $y$. As we derive $\mathbin{\tilde{x}}$ from a teacher LLM, this can be considered a form of symbolic knowledge distillation (West et al., 2022).
Dataset Class Train, Test Corpus Difficulty
@@
Polarity 2 $72\text{k}^{*},7.2\text{k}^{*}$ Products Medium
Category 23 $30\text{k}^{*},2.4\text{k}^{*}$ Products Medium
@@
IMDb 2 $20\text{k},25\text{k}$ Movies Medium
SST-2 2 $54\text{k},872$ Movies Medium
Table 1: Dataset statistics and our estimate of task inversion difficulty. Downsampled for convenience.